# Bibliografia.bib

### v1

@@ 5.1 Net Traded Value as a measure of market value
To quantify the financial impact of newly created tokens, we introduce the concept of NTV. This metric allows us to quantify the value these tokens, providing a clearer picture of the actual economic significance of new token creation. In particular, by distinguishing between honeypots and sellable tokens, NTV helps reveal that a substantial portion of market liquidity is, in reality, fictitious, as it remains trapped in honeypots. This distortion makes the market appear more valuable than it actually is, highlighting the discrepancy between perceived and real liquidity. Moreover, token NTV allows us to identify additional risks associated with newly created tokens, particularly two major pitfalls: the liquidation of a large portion of liquidity by the owner, leading to drastic price collapses, and rug pulls, where liquidity is removed from the pool. In the latter case, investors are left holding tokens that, despite appearing to have a certain market value, are effectively worthless due to the complete absence of liquidity, making them impossible to sell.
@@
where $v^{i}_{t}$ represents the signed token amount exchanged in the $t$-th swap for the $i$-th token, $N_{t}$ is the number of swaps considered, and $p^{i}(t=N_{t})$ denotes the token marginal price (in ETH) of the new coin at the time of the $N_{t}$-th swap for the $i$-th token. In Figure 4 we show the NTV of new tokens for $N_{t}=\{10,100,1000\}$ for both sellable and honeypot tokens.
(a) Histogram of NTV after 10 swaps. The minimum value for sellable tokens is -0.73 ETH while the maximum is 765.55 ETH. Instead for honeypots are respectively $-2.154\times 10^{15}$ ETH and $3.089\times 10^{13}$ ETH.
(b) Histogram of NTV after 100 swaps. The minimum value for sellable tokens is -22.06 ETH while the maximum is 1080.75 ETH. Instead for honeypots are respectively $-3.922\times 10^{13}$ ETH and $2.096\times 10^{13}$ ETH.
(c) Histogram of NTV after 1000 swaps. The minimum value for sellable tokens is -16.30 ETH while the maximum is 5787.21 ETH. Instead for honeypots are respectively -14.55 ETH and $1.331\times 10^{5}$ ETH.
@@
Figure 5: Scatter plot showing the maximum NTV value in the lifetime of all new tokens against the corresponding Liquidity value at that time. The $0.14\%$ of tokens with a maximum negative NTV are excluded.
@@
$N_{t}$ honeypot $\text{NTV}_{\text{tot}}$ (ETH) sellable $\text{NTV}_{\text{tot}}$ (ETH)
@@
Interesting insights can be extracted from analyzing extreme values of NTV, both negative and positive. Let us start with the negative ones. Indeed, as previously discussed, NTV incorporates the cumulative signed volume of recorded trades in its definition. When a token is created, the owner deposits an initial amount of liquidity into the pool, often using all or a portion of the total token supply. However, it is possible for the owner to retain a significant share of the token supply in their personal wallet. At any point, the owner may decide to sell these retained tokens in large quantities, causing the cumulative trade volume to turn negative and, in turn, to a negative NTV. After this action the price collapse, leaving investors with a token whose value has dramatically declined. The negative values primarily originate from honeypot tokens.
Action Timestamp Amount ETH Amount NT Reserve ETH Reserve NT Price (ETH)
@@ 5.2 Sandwich attacks
Using our strategy, only 0.28% of sellable tokens generate a profit exceeding 1 ETH. Excluding these outliers, the cumulative profit turns negative, indicating that success relies on a few tokens with extreme gains. This raises the question: what drives this small fraction of highly profitable trades? The answer lies in sandwich attacks.
A sandwich attack exploits price slippage by front-running a victim buy order. The attacker detects the victim trade in the mempool, submits a buy with a higher gas fee to execute first, inflating the price, and then sells immediately after the victim purchase. This process is optimized using FlashBots [flashbot2022], which bypasses the public mempool to prevent competition and ensure precise execution [MEV] by submitting transaction bundles. Uniswap AMM model and low-liquidity pools amplify these attacks, as even small buy orders can cause significant price slippage. Table 3 provides an example where a front-runner gains $s=0.0316$ ETH from a $25$ ETH investment. We estimate from our data that, among tokens with at least 100 swaps, approximately $68\%$ of honeypot tokens and $90\%$ of sellable tokens have experienced at least one sandwich attack event [^1] We define a sandwich attack event as one in which, at the same timestamp (within the same block), two swaps of opposite signs occur, where the absolute value of the swapped token amounts are similar within a $5\%$ threshold..
Our strategy takes advantage of some of these events, which are theoretically profitable when considering price evolution at swap time but not in reality. Exploiting these events generate the small percentage of extreme positive profits that, at least theoretically, make our strategy profitable overall. As sandwich attackers typically execute large trades relative to liquidity, the resulting price spike is substantial. By systematically capturing these price distortions, our approach remains profitable. Figure 6 illustrates the normalized price evolution over the first 350 swaps for all coins that yielded a profit greater than 0.5 ETH in our strategy. Sharp price peaks followed by immediate drops correspond to sandwich attacks, affecting both sellable tokens and honeypots. Our strategy exploits these peaks, securing profits before price reverts. Without these attacks, the strategy would yield negative returns and, thus, in real investments would not be exploitable to obtain positive profits.
@@
$\Delta x$ $\Delta y$ $x_{\rm new}$ $y_{\rm new}$ New Price Old Price Liquidity 25.$0$268464000 28.3199 35758537 7.920e-07 1.091e-08 31823 $0$62833 28.3699 35695704 7.948e-07 7.920e-07 31823 -25.0316 268463996 3.3382 304159700 1.098e-08 7.948e-07 31865
Table 3: Sandwich attack on the token AIDLE (pair address $=$ ”0xdab46ad9dc16d3de41fdcdc4eabd26fe8b188d91”) executed in the block 21560868. Here, New Price indicates the price after the swap, while Old Price the price before.
Figure 6: The figures show the normalized price trends using a min-max scaler for the first 350 swaps. The right panel displays all sellable coins that generated a profit greater than 0.5 ETH, while the left panel shows the same for honeypots. Sharp peaks reaching large values and immediately dropping can be observed for all coins, corresponding to sandwich attack events. There are only buy swap sandwiches because at the end of the three transactions, the attacker remains with only WETH.
Optimal Sandwich Strategy and its higher profitability in low-liquidity pools. Detecting pending transactions in the mempool of the Ethereum blockchain allows traders to anticipate the intentions of others before their transactions are executed. Specifically, by monitoring transactions targeting swap router contract addresses, it is possible to identify when a trader is about to buy a given quantity of a token, whose reserve is denoted with $y$, in exchange for a certain amount of Ethereum, whose reserve is $x$.
Determining the optimal investment amount for a sandwich attack is crucial [^2] In Ref.[heimbach2024sandwich], the authors derive the optimal strategy in the limit of huge liquidity or when the pool fee is zero ($r=1$).. Indeed, while increasing the investment might seem beneficial, in the limit of an infinitely large trade, the attacker would effectively perform a round-trip trade that results in losses due to pool fees and gas costs, without generating any real profit. Therefore, the optimal investment size must balance maximizing the price impact from the front-running trade while avoiding excessive slippage and transaction costs.
To analyze this in detail, we consider the execution order of the three key transactions involved in a sandwich attack:
$\Delta y_{a}=\frac{r\Delta x_{a}y}{x+r\Delta x_{a}}\,\,;$
where $r=1-f$ $(f=0.3\%\ll 1)$.
$\Delta y_{\epsilon}=\frac{r\Delta x_{\epsilon}(y-\frac{r\Delta x_{a}y}{x+ra})}{x+a+r\Delta x_{\epsilon}}\,\,;$
$s=\Delta x_{tot}-\Delta x_{a}\,\,.$
Note that in the case where the value of $\Delta x_{\epsilon}$ is 00 we obtain the amount $\Delta x_{\rm tot}$ characterized only by the pool fee which leads to an asymptotic loss given by
the maximum is for
@@
expressing $s_{\rm max}$ using the value of $\Delta x_{a}^{\rm max}\,$:
$s_{\rm max}=\frac{(\Delta x_{\epsilon}-fx)^{2}}{\Delta x_{\epsilon}}\,\,,$ (8)
@@
Figure 7: Visualization of the function $s$ in Eq.(5) as a function of the attacker input amount $\Delta x_{a}$. The plot is generated using the parameter values $x=1$, $\Delta x_{a}=1.3333$, $r=0.997$, and $\Delta x_{\epsilon}=0.1$. The red vertical line corresponds to the estimated maximum point of Equation (7).
For $\Delta x_{\epsilon}=0$, we recover the result previously derived in Equation (4), while for $r=1$, the maximum gain simplifies to $\Delta x_{\epsilon}$. A crucial insight from this analysis is that performing a sandwich attack in a highly liquid pool, i.e., with large $x$, leads to diminishing or even negative returns. This becomes clearer by taking the limit $x\rightarrow\infty$ in Equation (5), yielding:
$s(\Delta x_{a};\Delta x_{\epsilon})\rightarrow\Delta x_{a}(r^{2}-1)<0,\quad\text{for }x\rightarrow\infty.$ (10)
This result highlights a fundamental conclusion: sandwich attacks are far more profitable in low-liquidity pools. The reason is that in high-liquidity pools, the price impact of a single trade is much lower, making it harder for the attacker to extract significant profit from the victim trade. This insight may explain why sandwich attacks are particularly prevalent in newly created tokens, which typically have very low liquidity, making them ideal targets for such exploitative strategies.
When executing a trade on Uniswap V2, users can set a slippage tolerance, which defines the maximum acceptable price deviation from the expected execution price. This protects traders from excessive price impact caused by sudden market fluctuations or front-running attacks. In the context of sandwich attacks, this means that the victim effectively imposes an upper limit on the price they are willing to pay for the token. This corresponds to placing a horizontal line on the plot shown in Figure 7, intersecting the profit curve. If this line is positioned below the peak, the optimal strategy for the attacker is to invest precisely the amount that corresponds to the victim slippage tolerance. An interesting property of this scenario is the symmetry around the peak, implying that two different investment amounts—one smaller and one larger—can yield the same profit. On the other hand, if the slippage tolerance is higher than the peak, the attacker optimal profit is achieved at the peak itself.
@@ 5.3 The different behavior of new coins in swap and physical time
In the previous sections, the crucial distinction between swap time and physical time when analyzing price dynamics on Uniswap emerged. This differentiation is particularly relevant when evaluating the feasibility of profit opportunities, as real-world execution is constrained by block times and transaction ordering within the blockchain. In this section, we further investigate this aspect by studying the price evolution of newly created tokens during the initial phase of their existence. To achieve this, we analyze price trajectories using both swap time and physical time, applying clustering techniques [ClusterData] to identify recurring patterns and behaviors in the early life cycle of newly created tokens and providing a comparative view of how their prices evolve under these different temporal frameworks. By grouping tokens with similar price evolution, we can better understand the common trends of honeypots and sellable coins. Additionally, this clustering enables us to assess how different price behaviors influence potential profitability depending on whether trades are executed based on swap events or real-world block time constraints. As we will see, the findings reinforce the conclusions from the previous paragraphs regarding the limitations of our trading strategy. The discrepancies between swap time and physical time play a fundamental role in determining the actual accessibility of profit opportunities, as well as the risks associated with different trading approaches.
@@
Clustering in physical time. Let us now consider the case in which we consider price changes in physical time. We applied a dynamic time warping (DTW) method [Sakoe] on the financial data to categorize these tokens into clusters [Cobras]. This approach allowed us to group the sellable tokens and honeypot tokens based on their temporal price evolution (see Figure 10). By aligning and comparing the time series, the Dynamic Time Warping (DTW) facilitated the identification of distinct patterns and behaviors, enabling a deeper understanding of the dynamics within each cluster. We apply this algorithm to the time series of normalized prices, defined as the ratio between the difference of the price at a given time $t$ and its minimum, and the difference between its maximum and minimum. This normalization ensures that the price oscillates between 0, when it reaches its minimum value, and 1, when it attains its maximum.
@@
Figure 10: The figures show the mean and the one standard deviation bands of the trajectories resulting from the clustering, with $k=3$ clusters. We show the normalized price in function of the time expressed in seconds.
@@ 6 Conclusions
In this work, we carried on a study of new tokens on the Ethereum blockchain, addressing their security issues and their financial impact on Uniswap V2. With around 15 tokens created every hour, we explored the feasibility of implementing a trading strategy in this rapidly evolving market and analyzed the challenges associated with its execution. The primary motivation for investing in this sector lies in the high potential returns that these tokens can achieve within their short life cycle. To quantify their financial significance, we introduced the concept of NTV, which allows us to assess their actual economic value within the market. Investing in this market appears to be highly lucrative, as demonstrated by a simple buy & hold strategy. However, a closer examination of where the highest profits were generated reveals some intriguing insights.
First of all, the most profitable tokens were those associated with honeypots, which, by definition, prevent uninformed investors from selling their tokens once purchased. These honeypot tokens make up approximately 88% of the total, meaning that while they appear profitable on paper, they trap investors, making it impossible to realize gains.
@@
The remaining sellable tokens–which allow for resale–still yield a positive profit, even if with a lesser extent. This is because the strategy assumes an immediate sell after the buy swap in the first transaction of a sandwich attack. However, in practice, executing such precise timing is impossible since most of these attacks utilize FlashBots, bypassing the public mempool and preventing external traders from positioning their transactions accordingly.
We show that the high prevalence of sandwich attacks in newly launched cryptocurrencies is driven by their greater profitability in low-liquidity pools. To quantify the maximum extractable profit (excluding gas fees) from a sandwich attack, we derive an expression showing that the attacker profit is approximately equal to the victim trade size, minus adjustments that depend on the pool fee and the reserve size of WETH within the pair pool. In fact, deep liquidity pools mitigate sandwich attacks, as they minimize the attacker ability to move prices.
@@
The presence of sandwich attacks stresses the need to consider a trading strategy based on physical time rather than swap time. This distinction becomes evident when analyzing clustering patterns in both swap time and physical time. In swap-time clustering, price movements appear smoother and less dependent on transaction ordering. In physical-time clustering, distinct clusters emerge where price paths consistently increase in the first blocks. The most populated cluster in physical time shows that for the majority of tokens the price increases in the first block and it is immediately followed by a sharp decline, leading to the token death.
Investing in the first blocks of a new crypto is extremely difficult because we lack control over the precise position of our transaction within the block. MEV strategies, priority gas auctions, and sandwich attacks make it nearly impossible for retail traders to execute at optimal prices.
@@
Our study opens several avenues for future research, particularly in the areas of rug pulls, liquidity analysis, and early-block transaction dynamics. Future work should focus on developing more comprehensive methodologies to identify rug pulls at an earlier stage, potentially by analyzing smart contract characteristics before they are executed.
Acknowledgements

### v2

@@ 5.1 Net Traded Value as a measure of market value
To quantify the financial impact of newly created tokens [28], we introduce the concept of NTV. This metric allows us to quantify the value of these tokens, providing a clearer picture of the actual economic significance of new token creation. In particular, by distinguishing between honeypots and sellable tokens, NTV helps reveal that a substantial portion of market liquidity is, in reality, fictitious, as it remains trapped in honeypots. This distortion makes the market appear more valuable than it actually is, highlighting the discrepancy between perceived and real liquidity. Moreover, token NTV allows us to identify additional risks associated with newly created tokens, particularly two major pitfalls: the liquidation of a large portion of liquidity by the owner, leading to drastic price collapses, and rug pulls, where liquidity is removed from the pool. In the latter case, investors are left holding tokens that, despite appearing to have a certain market value, are effectively worthless due to the complete absence of liquidity, making them impossible to sell.
From an economic perspective, NTV is not intended as an absolute performance metric but as a proxy for the cumulative economic value extracted from a liquidity pool over time, expressed in WETH. By aggregating directional activity (inflows minus outflows) in value terms, NTV captures the net flow of economic value rather than raw traded volume, making it possible to distinguish between transient swap activity and persistent capital outflows linked to arbitrage or manipulation. Although our dataset does not include all pools or Uniswap versions, NTV remains internally consistent at the pool level, since MEV and arbitrage cycles are typically local and self-contained. Furthermore, as discussed earlier, the presence of parallel pools for tokens originally created on Uniswap V2 across other DEXes is negligible, and thus does not materially affect the interpretation of NTV. Consequently, relative variations in NTV provide an indication of trading pressure and value extraction dynamics within each observed market.
@@
where $v^{i}_{t}$ represents the signed token amount exchanged in the $t$-th swap for the $i$-th token, $N_{t}$ is the number of swaps considered, and $p^{i}(t=N_{t})$ denotes the token marginal price (in WETH) of the new coin at the time of the $N_{t}$-th swap for the $i$-th token.
In Figure 3 we show the distribution of NTV of new tokens for $N_{t}=\{100,1000\}$ for both sellable and honeypot tokens.
(a) Histogram of NTV after 100 swaps. The minimum value for sellable tokens is -22.06 WETH while the maximum is 1080.75 WETH. Instead for honeypots are respectively $3.922\times 10^{13}$ WETH and $2.096\times 10^{13}$ WETH.
(b) Histogram of NTV after 1000 swaps. The minimum value for sellable tokens is -16.30 WETH while the maximum is 5787.21 WETH. Instead for honeypots are respectively -14.55 WETH and $1.331\times 10^{5}$ WETH.
@@
Figure 4: Scatter plot showing the maximum NTV value in the lifetime of all new tokens against the corresponding liquidity value at that time. The $0.14\%$ of tokens with a maximum negative NTV are excluded. A negative NTV occurs when a hidden supply held by the token creator, which was not included in the initial mint, is used to sell into the pool.
@@
$N_{t}$ honeypot $\text{NTV}_{\text{tot}}$ (WETH) sellable $\text{NTV}_{\text{tot}}$ (WETH)
@@
Interesting insights can be extracted from analyzing extreme values of NTV, both negative and positive. Let us start with the negative ones. Indeed, as previously discussed, NTV incorporates the cumulative signed volume of recorded trades in its definition. When a token is created, the owner deposits an initial amount of liquidity into the pool, often using all or a portion of the total token supply. However, it is possible for the owner to retain a significant share of the token supply in their personal wallet. At any point, the owner may decide to sell these retained tokens in large quantities, causing the cumulative trade volume to turn negative and, in turn, to obtain a negative NTV. After this action the price collapses, leaving investors with a token whose value has dramatically declined. The negative values primarily originate from honeypot tokens.
Action Timestamp Amount WETH Amount NT Reserve WETH Reserve NT Price (WETH)
@@ 5.2 Sandwich attacks
Using our strategy, only 0.28% of sellable tokens generate a profit exceeding 1 WETH. Excluding these outliers, the cumulative profit turns negative, indicating that success relies on a few tokens with extreme gains. This raises the question: what drives this small fraction of highly profitable trades? The answer lies in sandwich attacks.
A sandwich attack exploits price slippage by front-running a victim buy order. The attacker detects the victim trade in the mempool, submits a buy with a higher gas fee to execute first, inflating the price, and then sells immediately after the victim purchase. This process is optimized using FlashBots [29], which bypasses the public mempool to prevent competition and ensure precise execution [30, 31] by submitting transaction bundles. Uniswap AMM model and low-liquidity pools amplify these attacks, as even small buy orders can cause significant price slippage. We estimate from our data that, among tokens with at least 100 swaps, approximately $68\%$ of honeypot tokens and $90\%$ of sellable tokens have experienced at least one sandwich attack event [^1] We define a sandwich attack event as one in which, at the same timestamp (within the same block), two swaps of opposite signs occur, where the absolute value of the swapped token amounts are similar within a $5\%$ threshold..
Our strategy takes advantage of some of these events, which are theoretically profitable when considering price evolution at swap time but not in reality. Exploiting these events generate the small percentage of extreme positive profits that, at least theoretically, make our strategy profitable overall. As sandwich attackers typically execute large trades relative to liquidity, the resulting price spike is substantial. By systematically capturing these price distortions, our approach remains profitable. Sharp price peaks followed by immediate drops correspond to sandwich attacks, affecting both sellable tokens and honeypots. Our strategy exploits these peaks, securing profits before price reverts. Without these attacks, the strategy would yield negative returns and, thus, in real investments would not be exploitable to obtain positive profits.
@@
Optimal Sandwich Strategy and its higher profitability in low-liquidity pools. Determining the optimal investment amount for a sandwich attack is crucial [^2] In Ref.[32], the authors derive the optimal strategy in the limit of huge liquidity or when the pool fee is zero ($r=1$).. Indeed, while increasing the investment might seem beneficial, in the limit of an infinitely large trade, the attacker would effectively perform a round-trip trade that results in losses due to pool fees and gas costs, without generating any real profit. Therefore, the optimal investment size must balance maximizing the price impact from the front-running trade while avoiding excessive slippage and transaction costs.
Introducing $s$ as the gain in WETH for the sandwich attacker, we define the maximum value $s_{\rm max}$ as occurring when the amount of the WETH swapped by the attacker in the first transaction, $\Delta x_{a}^{\rm max}$ , is (see Appendix C for details):
@@
where $\Delta x_{\epsilon}$ is the amount of Ethereum sold by the the victim and $x$ is the Ethereum reserve into the pool. Expressing $s_{\rm max}$ using the value of $\Delta x_{a}^{\rm max}\,$:
$s_{\rm max}=\frac{(\Delta x_{\epsilon}-fx)^{2}}{\Delta x_{\epsilon}}\,\,.$ (5)
@@
A crucial insight from this analysis is that performing a sandwich attack in a highly liquid pool, i.e., with large $x$, leads to diminishing or even negative returns. This becomes clearer by taking the limit $x\rightarrow\infty$ in Equation (9), yielding:
$\centering s(\Delta x_{a};\Delta x_{\epsilon})\rightarrow\Delta x_{a}(r^{2}-1)<0,\quad\text{for }x\rightarrow\infty.\@add@centering$ (6)
This result stresses a fundamental conclusion: sandwich attacks are far more profitable in low-liquidity pools. The reason is that in high-liquidity pools, the price impact of a single trade is much lower, making it harder for the attacker to extract significant profit from the victim trade. This insight may explain why sandwich attacks are particularly prevalent in newly created tokens, which typically have very low liquidity, making them ideal targets for such exploitative strategies.
When executing a trade on Uniswap V2, users can set a slippage tolerance, i.e. $\eta\in[0,1]$, which defines the maximum acceptable price deviation from the expected execution price. This protects traders from excessive price impact caused by sudden market fluctuations or front-running attacks. In the context of sandwich attacks, this means that the victim effectively imposes an upper limit on the price they are willing to pay for the token. This corresponds to placing a vertical line on the plot shown in Figure 15, intersecting the profit curve. If this line is positioned below the peak, the optimal strategy for the attacker is to invest precisely the amount that corresponds to the victim slippage tolerance. An interesting property of this scenario is the symmetry around the peak, implying that two different investment amounts—one smaller and one larger—can yield the same profit. On the other hand, if the slippage tolerance is higher than the peak, the attacker optimal profit is achieved at the peak itself.
@@ 5.3 The different behavior of new coins in swap and physical time
In the previous sections, the crucial distinction between swap time and physical time when analyzing price dynamics on Uniswap emerged. This differentiation is particularly relevant when evaluating the feasibility of profit opportunities, as real-world execution is constrained by block times and transaction ordering within the blockchain. In this section, we further investigate this aspect by studying the price evolution of newly created tokens during the initial phase of their existence. To achieve this, we analyze price trajectories using both swap time and physical time, applying clustering techniques [33] to identify recurring patterns and behaviors in the early life cycle of newly created tokens and providing a comparative view of how their prices evolve under these different temporal frameworks. By grouping tokens with similar price evolution, we can better understand the common trends of honeypots and sellable coins. Additionally, this clustering enables us to assess how different price behaviors influence potential profitability depending on whether trades are executed based on swap events or real-world block time constraints. As we will see, the findings reinforce the conclusions from the previous paragraphs regarding the limitations of our trading strategy. The discrepancies between swap time and physical time play a fundamental role in determining the actual accessibility of profit opportunities, as well as the risks associated with different trading approaches. A detailed discussion of the clustering results, including the elbow curve and PCA visualizations for $k=2$ and $k=3$ clusters, is provided in Appendix D.
@@
Clustering in physical time. Let us now consider the case in which we consider price changes in physical time. We applied a dynamic time warping (DTW) method [34] on the financial data to categorize these tokens into clusters [35]. This approach allowed us to group the sellable tokens and honeypot tokens based on their temporal price evolution (see Figure 7). By aligning and comparing the time series, the Dynamic Time Warping (DTW) facilitated the identification of distinct patterns and behaviors, enabling a deeper understanding of the dynamics within each cluster. We apply this algorithm to the time series of normalized prices. This normalization ensures that the price oscillates between 0, when it reaches its minimum value, and 1, when it attains its maximum.
@@
Figure 7: The figures show the mean and the one standard deviation bands of the trajectories resulting from the clustering, with $k=3$ clusters (see App. D for the elbow curve analysis). We show the normalized price in function of the time expressed in seconds.
@@ 6 Conclusions
In this work, we carried on a study of new tokens on the Ethereum blockchain, addressing their security issues and their financial impact on Uniswap V2. With around 15 tokens created every hour, we explored the feasibility of implementing a trading strategy in this rapidly evolving market and analyzed the challenges associated with its execution. The primary motivation for investing in this sector lies in the high potential returns that these tokens can achieve within their short life cycle. To quantify their financial significance, we introduced the concept of NTV, which allows us to assess their actual economic value within the market. Investing in this market appears to be highly lucrative, as demonstrated by a simple buy & hold strategy. However, a closer examination of where the highest profits were generated reveals some intriguing insights. First of all, the most profitable tokens were those associated with honeypots, which, by definition, prevent uninformed investors from selling their tokens once purchased. These honeypot tokens make up approximately 88% of the total, meaning that while they appear profitable on paper, they trap investors, making it impossible to realize gains.
@@
The remaining sellable tokens–which allow for resale–still yield a positive profit, even if to a lesser extent. This is because the strategy assumes an immediate sell after the buy swap in the first transaction of a sandwich attack. However, in practice, executing such precise timing is impossible since most of these attacks utilize FlashBots, bypassing the public mempool and preventing external traders from positioning their transactions accordingly.We show that the high prevalence of sandwich attacks in newly launched cryptocurrencies is driven by their greater profitability in low-liquidity pools. To quantify the maximum extractable profit (excluding gas fees) from a sandwich attack, we derive an expression showing that the attacker profit is approximately equal to the victim trade size, minus adjustments that depend on the pool fee and the reserve size of WETH within the pair pool. In fact, deep liquidity pools mitigate sandwich attacks, as they minimize the attacker ability to move prices.
@@
The presence of sandwich attacks stresses the need to consider a trading strategy based on physical time rather than swap time. This distinction becomes evident when analyzing clustering patterns in both swap time and physical time. In swap-time clustering, price movements appear smoother and less dependent on transaction ordering. In physical-time clustering, distinct clusters emerge where price paths consistently increase in the first blocks. The most populated cluster in physical time shows that for the majority of tokens the price increases in the first block and it is immediately followed by a sharp decline, leading to the token death [36]. Investing in the first blocks of a new crypto is extremely difficult because we lack control over the precise position of our transaction within the block. MEV strategies, priority gas auctions, and sandwich attacks make it nearly impossible for retail traders to execute at optimal prices.
As a first step, prospective investors approaching newly deployed tokens should rely on publicly accessible security platforms such as honeypot.is or GoPlus, which are free and provide preliminary risk assessments. However, these tools may not always offer complete coverage, particularly during the first few seconds of a token lifecycle. Certain quantitative properties can raise immediate red flags of potential honeypot behavior. For example, a small number of swaps involving unusually high amounts of WETH may indicate an artificial price increase designed to attract buyers. Similarly, a high NTV value combined with low liquidity often signals a possible rug pull, while a high NTV alone may suggest market manipulation. Most of this information is readily available through online token-scanning platforms, where the NTV can be seen as a surrogate for the market capitalization commonly displayed on such websites. For more in-depth and real-time assessments immediately after a token creation, investors may also rely on paid security services, which act as a safer bridge between uninformed users and professional researchers and/or smart contract developers.
@@
Moreover, our study opens several avenues for future research, particularly in the areas of rug pulls, liquidity analysis, and early-block transaction dynamics. Future work should focus on developing more comprehensive methodologies to identify rug pulls at an earlier stage, potentially by analyzing smart contract characteristics before they are executed.
Abbreviations
  • DEX = Decentralized EXchange;
  • V2 = Version 2;
  • AMM = Automated Market Maker;
  • ETH = Ethereum;
  • WETH = Wrapped Ethereum;
  • NTV = Net Traded Value;
  • DTW = Dynamic Time Warping;
  • MEV = Maximal Extractable Value;
Declarations
1. Ethics approval and consent to participate
Not applicable.
2. Consent for publication
The authors consent to the publication of their article in the journal EPJ Data Science.
3. Availability of data and material
The data are publicly available through interaction with the Ethereum blockchain.
4. Competing interests
The authors declare no competing interests.
5. Funding
This research received funding from the following grants: